y+4>= ___
x+2=7 ___
4m=12 ___
3r+6 ___
x+1<6 ___
x-2>5 ___
x+9=4 ___
w/3=2 ___
x-1=4 ___
n+3>5 ___
9x-3=6 ___

Answer by College Student(505)

y+4>= y>=-4
x+2=7 x=5
4m=12 m=12/4 ...simplified... m=3
3r+6 <-- stays as is
x+1<6 x<5
x-2>5 x>7
x+9=4 x=-5
w/3=2 w=2/3
x-1=4 x=5
n+3>5 n>5-3 ...simplified... n>2
9x-3=6 9x=6+3 ...same as... 9x=9 ...simplified... x=9/9 ...same as... x=1